# Symmetry aspects in an ideal Bose gas at thermal equilibrium

### v1

@@ Abstract
The theory of ideal gases is supplemented by a numerical quantum field description with a two-dimensional non-local order parameter that allows the modeling of wave-like atomic correlations and interference effects in the limit of low atomic densities. From the present model, it is possible to explain certain new fundamental symmetry aspects of ideal and very weakly interacting Bose gases, like the forward propagation of time and the relation to the breaking and preservation of phase gauge symmetry in solids. In the present formalism, the propagation of one-directional time arises from the pre-defined and equivalent convergence of independent quantum fields towards the Boltzmann equilibrium, and it is shown that Glauber coherent states are related to the definition of the quantized field. As a direct consequence of thermal equilibrium, the present model describes spontaneous symmetry breaking, which is originally only known from the definition of a specific gauge field from Elizur’s theorem for local gauge fields, as a continuous physical rather than a purely formal mathematical process.
@@ I Introduction
Theoretically, the theory of ideal gases provides a straightforward tool to understand and model the dynamics of many particles for the entire range below and above the critical temperature for Bose-Einstein condensation ref-6 ; ref-7 . Such experimental realizations of Bose-Einstein condensates with minimal interaction strength (described within first-order perturbation theory by the s-wave scattering length) are typically well-covered by the model results of kinetic gas theories, i.e. in the limit of low densities and high temperatures at zero interaction strength ref-8 . Yet the fundamental results and applicational practices of ideal gas theory are well established, so far, more needs to be known about simple theoretical formulations to explain the relation of reversible microscopic many-particle dynamics and the time direction of many-body quantum systems ref-9 ; ref-10
@@
In the context of ideal gases, irreversibility is understood as entropy maximization at the Boltzmann equilibrium, as formulated by the second law of thermodynamics. A daily observation of experiments is that time passes irreversibly in one direction, which can e. g. be counted with an atomic clock that measures the pings of oscillating photons or atoms, respectively, at the edges of a microwave cavity or by the number of decays of atoms in an atomic crystal, to define a unit of time. However, since non-interacting gases do not obey dissipation or inelastic collisions, the dynamics of ideal many-body processes are completely reversible in time, and the fundamental process for breaking time or phase gauge symmetries must not depend on the interactions of a many-particle system ref-11 . It is, therefore, in particular, not trivial to combine the theory of ideal gases with the concept of forward propagating time - since, as argued before, time principally obeys a (reflection) symmetry for non-interacting atomic systems.
@@
Relating ideal gases to the concept of spontaneous symmetry breaking in Bose-Einstein condensates in terms of breaking the phase gauge invariance in the quantum system is another delicate task ref-12 , and one may further ask whether or if the forward propagation of a time variable as defined in physical units is it related to a sort of broken phase gauge symmetry in a non-interacting many-body quantum system? What is the relation of one-directional time propagation, symmetry breaking at the phase transition for Bose-Einstein condensation, and the occurrence of a quantum field with a well-defined, but presumably random phase that describes the wave nature of the quantum system?
@@
To answer these questions that are currently not straightforwardly explained by the standard theory of ideal gases, the present work addresses and summarizes the fundamental results of a number-conserving quantum field theory that can be applied to Bose gases in the limit of vanishing interactions. First, the symmetry preserving unitary time evolution according to the ideal gas theory is introduced, before the extensions to the description of an ideal gas in terms of a non-local order parameter in complex number representation is explained. In the framework of this non-local quantum field, theoretical description for a total number of $N$ particles at a definite temperature $T$, decay, and oscillation rates associated with partial chemical potentials can be numerically derived in complex number representation. This quantum model, in particular, explains the one-directional propagation of time with an average zero phase of the quantum wave field from the interference of ideally disjoint forward and backward propagating wave fields at the Boltzmann equilibrium ref-9 with vanishing total chemical potential.
@@
Finally, this explains that spontaneously broken gauge symmetries of the gauge fields for the Schrödinger equation occur ab initio as a physical process from the thermodynamical constraints of the quantum system for a theory with non-local order parameters, without the requirements of formally defining a certain gauge for the quantum field, but only by implementing the physical boundary condition of a Boltzmann thermal equilibrium into the theoretical model.
@@ II.1 Unitary time evolution
$\hat{U}(t)={\rm exp}\left[-\frac{i\hat{H}t}{\hbar}\right]\ ,$ (1)
where $\hat{H}=\sum_{\epsilon_{\bf k}}\epsilon_{\bf k}\hat{a}^{\dagger}_{\bf k}\hat{a}_{\bf k}$ represents the energy operator of the quantum many-particle system in second quantization. In Eq. (1), the variables $\epsilon_{\bf k}$ define the energy of a single particle in the external trapping potential, and the operators $\hat{a}^{\dagger}_{\bf k}$ and $\hat{a}_{\bf k}$ describe the creation and destruction of quantum states of a single particle that populates a certain state with the corresponding energy $\epsilon_{\bf k}$. From the eigenvalue equation
@@ II.2 Quantum field and order parameter
that describes realizations of the many-body quantum states of $N$ non-interacting particles with random coefficients $b_{\bf k}$. From Elizur’s theorem ref-16 , one may learn that there exist no spontaneously broken local gauge symmetries without fixing the field gauge, i.e., since the phase gauge symmetry of quantum systems is broken at the phase transition for Bose-Einstein condensation, it is formally motivated to define a non-local spatially averaged order parameter for the Bose gas that describes phase transitions and the corresponding symmetry breaking of the gauge fields globally with a single numerical value ref-17 .
@@ II.3 Elementary model for time evolution
It is only at the Boltzmann equilibrium with vanishing chemical potential (which is derived for the equilibrium with the thermal external environment from particle number conservation at definite temperature), where both fields are exactly equal at multiples of this time interval, where the total system interacts with the thermal environment, which can be regarded as a measurement process that defines a macroscopic variable with the temperature $T$. Thus, the two field components $\psi_{+}(t)$ and $\psi_{-}(t)$ entail a symmetry property concerning a period $\tau_{0}$ that leads to the (parameter dependent) quantization of the time variable, $t=m\tau_{0}$, with integer values of $m$, where
@@ II.4 Representations with Glauber coherent states
A very intuitive and well-established representation of photonic eigenstates of the annihilation operator is defined by the Glauber coherent quantum states.
@@
which in the present formalism can be related to the quantum field of the ideal Bose gas at thermal equilibrium with the representation of the field in the complex number plane $\alpha=\psi=|\psi|{\rm e}^{i\phi}$ and the distribution $P(\alpha)$ which leads to
@@
$i\hbar\frac{\partial|\psi\rangle}{\partial t}=\hat{H}|\psi\rangle\ ,$ (12)
@@
$|\psi\rangle={\rm exp}\left[\frac{-i\hat{H}t}{\hbar}\right]|\psi_{0}\rangle$ (13)
@@
Figure 1: (Color online) Monte-Carlo simulation of the distribution $P(\alpha)$ with $5\times 10^{6}$ samples of the wave fields as defined in Eq. (7) to numerically model the quantum field distribution in units of Glauber coherent quantum states. Parameters for the numerical simulation are set to $\omega_{x}=2\pi\times 125.0$ Hz, $\omega_{y}=2\pi\times 75.0$ Hz and $\omega_{z}=2\pi\times 25.0$ Hz at $T=15.0$ nK with $N=5000$ atoms and few hundred considered eigenmodes.
Figure 2: (Color online) Figures highlight the broken phase gauge symmetry in the complex plane. The upper left figure shows the distribution of the condensate chemical potential at thermal equilibrium in the complex plane, which is composed of a ring with a constant radius in the complex plane and a symmetry-breaking constant part in the direction of the real axis. Upper right and lower left figures show the distribution of the imaginary parts (upper right) and the real parts (lower left) of the condensate wave field obtained from Eqs. (7), (9) und (17). The lower right figure explains the broken symmetry of the wave field’s aggregate phases, i.e., condensate and non-condensate parts (left-shifted and right-shifted circles in the representation of numbers in the complex plane). Parameters for the numerical simulation are set to $\omega_{x}=2\pi\times 125.0$ Hz, $\omega_{y}=2\pi\times 75.0$ Hz and $\omega_{z}=2\pi\times 25.0$ Hz at $T=25.0$ nK.
@@ II.5 Correlations of the relative phase
with the parameter $\chi(\Delta\phi,\beta)$ defined as the coherent sum of forward and backward propagating quantum fields $\chi(\Delta\phi,\beta)=\psi_{+}(\beta)+{\rm e}^{i\Delta\phi}\psi_{-}(\beta)$. An approximate analytical formula for the correlation function $c(\chi^{\star},\chi)$ is given by a Gaussian distribution around the average zero phase of $\chi(\Delta\phi=0,\beta)$, i.e.
$c(\chi^{\star},\chi)=c(\Delta\phi)=\frac{{\rm exp}\left[-\Delta\phi^{2}\right]}{\mathcal{N}}\ ,$ (15)
@@
More sophisticated numerical analysis shows mathematically more complex correlation functions between forward and backward propagating wave fields with zero average phases, which is shown in the quantitative analysis of the next chapter III. The Schrödinger equation in Eq. (12) once more indicates that spontaneous symmetry breaking isn’t related to the occurrence of a random, but to a pre-defined phase of the quantum field, since the physical constraints in a Bose gas intrinsically pick out a specific zero phase at the Boltzmann equilibrium, while the total distribution of all possible gauge fields is still symmetry preserving.
@@ II.6 Spontaneous phase gauge symmetry breaking
The concept of spontaneous phase gauge symmetry-breaking explains a certain sudden change in the symmetry properties of the gauge fields that describe a solid below the critical temperature. As an example, one may consider a magnet that changes its magnetization from average zero to a well-defined direction with components of spin up and spin down in the z-direction of the magnet below a specific Curie temperature for a characteristic phase transition. While the experimental observation of symmetry breaking is understood straightforwardly, the formulation of a consistent theory that explains spontaneous symmetry breaking as a physical rather than a purely mathematical process so far hasn’t been consistently formulated.
Figure 3: (Color online) Figure shows the distribution of the condensate wave field with the constraint of the Boltzmann equilibrium at the same experimental conditions as in Fig. 2. Instead of the continuously distributed wave field in the complex plane, the wave field shows a star-shaped linear spreading from zero to a finite value in the direction of quantized angles $\phi$.
Any initial state of the formal solution in Eq. (13) with a specific initial phase defined by $\psi_{0}=|\psi|{\rm e}^{i\phi_{0}}$ defines a separate and non-unique solution of the Schrödinger equation. The corresponding gauge fields are symmetric as long as no physical constraint is assumed for the quantum field’s chemical potential and are defined by Glauber coherent states that all preserve the symmetry of the Schrödinger equation. Considering only manifolds of the Boltzmann equilibrium with vanishing chemical potential $\mu=0$ leads to spontaneous symmetry breaking of the gauge fields as a consequence of a physical constraint, the Boltzmann equilibrium, i.e., to the distribution $P(\alpha)=\delta(\alpha-\alpha_{0})$, without the requirement of formally defining a certain symmetry-breaking gauge for the wave field. The projection of the total wave field onto the manifold of the Boltzmann equilibrium finally also implies the propagation of a one-directional time, i. e. ”the forward direction of time”. From Eqs. (6) - (8), it is thus likely to understand that physical time passes in only one direction, since at the Boltzmann equilibrium, where disjoint forward and backward propagating components of the quantum field are effectively mixed and equal, the partial waves interfere and lead to a well-defined mathematical form of the quantum field that doesn’t diverge numerically also below the critical temperature for Bose-Einstein condensation.
@@
Considering a quantum field without local representations of the particle’s wave functions facilitates the formulation of symmetry breaking between the different components (aggregate phases) of the quantum system. As seen before, in mathematical terms, the concept of macroscopic symmetry breaking in ideal Bose gases can be understood by the introduction of a non-local order parameter as defined by
@@
within the framework of the present number-conserving quantum field theory. Since a total number of $N$ non-interacting atoms is assumed, the total average field $\psi=0$ (order parameter of the total field) vanishes. This is, in particular, described and accounted for by Eq. (9) which represents the conservation of the total number of particles. From this assumption, it is relatively straightforward to recognize that the condensate and non-condensate aggregate phases of the Bose gas are separated by an angle of $\phi=\pi$, i.e.
@@
hence, interpreting the spontaneously broken gauge symmetry of the order parameter below the critical temperature is also rather straightforward in terms of the fields in Eq. (17).

### v2

# Physical aspects of symmetry breaking in an ideal Bose gas at thermal equilibrium

@@ Abstract
The theory of ideal gases is supplemented by a numerical quantum field description with a two-dimensional non-local order parameter that allows the modeling of wave-like atomic correlations and interference effects in the limit of low atomic densities. From the present model, it is possible to explain certain new fundamental symmetry aspects of ideal and very weakly interacting Bose gases, like the forward propagation of time and the relation to the breaking and preservation of phase gauge symmetry in solids. In the present formalism, the propagation of one-directional time arises from the pre-defined and equivalent convergence of independent quantum fields towards the Boltzmann equilibrium, and it is shown that Glauber coherent states are related to the definition of the quantized field. Coherently coupling condensate and non-condensate parts as a direct consequence of the increasing quantum coherence time between the different quantum field components in the Bose gas from cooling to below the critical temperature, the present model describes symmetry breaking, which is originally known from the definition of a specific gauge field from Elitzur’s theorem for local gauge fields, as a global physical rather than a purely formal mathematical process.
@@ I Introduction
Theoretically, the theory of ideal gases provides a straightforward tool to understand and model the dynamics of many particles for the entire range below and above the critical temperature for Bose-Einstein condensation ref-6 ; ref-7 . Such experimental realizations of Bose-Einstein condensates with minimal interaction strength (described within first-order perturbation theory by the s-wave scattering length) are typically well-covered by the model results of kinetic gas theories, i.e. in the limit of low densities and high temperatures at zero interaction strength ref-8 . Yet the fundamental results and applicational practices of ideal gas theory are well established, so far, more needs to be known about simple theoretical formulations to explain the relation of reversible microscopic many-particle dynamics and the time direction of many-body quantum systems ref-9 ; ref-10 .
@@
In the context of ideal gases, irreversibility is understood as entropy maximization at the Boltzmann equilibrium, as formulated by the second law of thermodynamics. A daily observation of experiments is that time passes irreversibly in one direction, which can e. g. be counted with an atomic clock that measures the pings of oscillating photons or atoms, respectively, at the edges of a microwave cavity or by the number of decays of atoms in an atomic crystal, to define a unit of time. However, since non-interacting gases do not obey dissipation or inelastic collisions, the dynamics of ideal many-body processes are completely reversible in time, and the fundamental process for breaking time or phase gauge symmetries must not depend on the interactions of a many-particle system ref-11 .
@@
It is, therefore, in particular, not trivial to combine the theory of ideal gases with the concept of forward propagating time - since, as argued before, time principally obeys a (reflection) symmetry for non-interacting atomic systems. Relating ideal gases to the concept of spontaneous symmetry breaking in Bose-Einstein condensates in terms of breaking the phase gauge invariance in the quantum system is another delicate task ref-12 , and one may further ask whether or if the forward propagation of a time variable as defined in physical units is related to a sort of broken phase gauge symmetry in a non-interacting many-body quantum system? What is the relation of one-directional time propagation, symmetry breaking at the phase transition for Bose-Einstein condensation, and the occurrence of a quantum field with a well-defined, but presumably random phase that describes the wave nature of the quantum system?
@@
To answer these questions that are currently not straightforwardly explained by the standard theory of ideal gases, the present work explains the context of the fundamental results of a number-conserving quantum field theory to representations of Bose-Einstein condensates in terms of Glauber coherent states. First, the symmetry preserving unitary time evolution according to the ideal gas theory is introduced, before extensions to the description of an ideal gas in terms of a non-local order parameter in complex number representation is explained. In the framework of this non-local quantum field theory for a total number of $N$ particles at a definite temperature $T$, decay, and oscillation rates associated with partial chemical potentials can be numerically derived in complex number representation. This quantum model, in particular, explains the one-directional propagation of time with an average zero phase of the quantum wave field from the interference of ideally disjoint forward and backward propagating wave fields at the Boltzmann equilibrium ref-9 with vanishing total chemical potential. The scaling of the coherence time indicates that the total quantum field is composed of a joint combination of two field components (a field generated from the condensate mode population and a second field created from the rest of the single particle modes) leading to a symmetric total field distribution that builds an extended ring in the complex plane above the critical temperature. This field decays into a point-like field of zero average value as a consequence of the increasing coherence time between condensate and non-condensate subspaces below the critical temperature when the coherence time approaches values of the minimal time uncertainty.
@@
Finally, this explains that forward time propagation and spontaneously broken gauge symmetries of the Schrödinger equation’s gauge fields occur as a physical process below the critical temperature, originating from quantum coherence of the quantum states in a theory with non-local order parameters that gauges the average total field to zero and further implements the physical boundary condition of a Boltzmann thermal equilibrium into the theoretical model.
@@ II.1 Unitary time evolution
$\hat{U}(t)={\rm exp}\left[-\frac{i\hat{\mathcal{H}}t}{\hbar}\right]\ ,$ (1)
where ^$\hat{\mathcal{H}}=\sum_{\epsilon_{\bf k}}\epsilon_{\bf k}\hat{a}^{\dagger}_{\bf k}\hat{a}_{\bf k}$ represents the energy operator of the quantum many-particle system in second quantization. In Eq. (1), the variables $\epsilon_{\bf k}$ define the energy of a single particle in the external trapping potential, and the operators $\hat{a}^{\dagger}_{\bf k}$ and $\hat{a}_{\bf k}$ describe the creation and destruction of quantum states of a single particle that populates a certain state with the corresponding energy $\epsilon_{\bf k}$. From the eigenvalue equation
@@ II.2 Quantum field and order parameter
that describes realizations of the many-body quantum states of $N$ non-interacting particles with random coefficients $b_{\bf k}$. From Elitzur’s theorem, ref-16 , one may learn that there exist no spontaneously broken local gauge symmetries without fixing the field gauge, i.e., since the phase gauge symmetry of three-dimensional quantum systems is suspected to be broken at the phase transition for Bose-Einstein condensation, it is formally motivated to define a non-local spatially averaged order parameter for the Bose gas that describes phase transitions and the corresponding symmetry breaking of the gauge fields globally with a single numerical value ref-17 .
@@ II.3 Elementary model for time evolution
It is only at the Boltzmann equilibrium with vanishing chemical potential (which is derived for the equilibrium with the thermal external environment from particle number conservation at definite temperature), where both fields are exactly equal at multiples of this time interval and the total system interacts with the thermal environment. This can be regarded as a measurement process that defines a macroscopic variable with the temperature $T$. Thus, the two field components $\psi_{+}(t)$ and $\psi_{-}(t)$ entail a symmetry property concerning a period $\tau_{0}$ that leads to the (parameter dependent) quantization of the time variable, $t=m\tau_{0}$, with integer values of $m$, where
@@ II.4 Representations with Glauber coherent states
A very intuitive and well-established representation of photonic eigenstates of the annihilation operator is defined by the Glauber coherent quantum states
@@
which in the present formalism can be related to the quantum field of the ideal Bose gas at thermal equilibrium with the representation of the field in the complex number plane $\alpha=\sqrt{N}\psi=\sqrt{N}|\psi|{\rm e}^{i\phi}$ and the distribution $P(\alpha)$ which leads to
@@
$i\hbar\frac{\partial|\psi\rangle}{\partial t}=\hat{\mathcal{H}}|\psi\rangle\ ,$ (12)
@@
$|\psi\rangle={\rm exp}\left[\frac{-i\hat{\mathcal{H}}t}{\hbar}\right]|\psi_{0}\rangle$ (13)
@@
Figure 1: (Color online) Monte-Carlo simulation of the distribution $P(\alpha)$ with $5\times 10^{6}$ samples of the (combined) wave fields as defined in Eq. (7) to numerically model the quantum field distribution in units of Glauber coherent quantum states. Parameters for the numerical simulation are set to $\omega_{x}=2\pi\times 125.0$ Hz, $\omega_{y}=2\pi\times 75.0$ Hz and $\omega_{z}=2\pi\times 25.0$ Hz at $T=15.0$ nK with $N=5000$ atoms and few hundred considered eigenmodes.
Figure 2: (Color online) Figures highlight the broken phase gauge symmetry in the complex plane. The upper left figure shows the distribution of the condensate chemical potential at thermal equilibrium in the complex plane, which is composed of a ring with a constant radius in the complex plane and a symmetry-breaking constant part in the direction of the real axis. Upper right and lower left figures show the distribution of the imaginary parts (upper right) and the real parts (lower left) of the condensate wave field obtained from Eqs. (7), (9) und (17). The lower right figure explains the broken symmetry of the wave field’s aggregate phases, i.e., condensate and non-condensate parts (left-shifted and right-shifted circles in the representation of numbers in the complex plane). Parameters for the numerical simulation are set to $\omega_{x}=2\pi\times 125.0$ Hz, $\omega_{y}=2\pi\times 75.0$ Hz and $\omega_{z}=2\pi\times 25.0$ Hz at $T=25.0$ nK with $N=5000$ atoms.
@@ II.5 Correlations of the relative phase
with the parameter $\chi(\Delta\phi,\beta)$ defined as the coherent sum of forward and backward propagating quantum fields $\chi(\Delta\phi,\beta)=\psi_{+}(\beta)+{\rm e}^{i\Delta\phi}\psi_{-}(\beta)$. An approximate analytical formula for the correlation function $c(\chi^{\star},\chi)$ is given by a Gaussian distribution around the average zero phases of $\chi(\Delta\overline{\phi}=0,\beta)$, i.e.
$c(\chi^{\star},\chi)=c(\Delta\phi)\approx\frac{{\rm exp}\left[-\Delta\phi^{2}\right]}{\mathcal{N}}\ ,$ (15)
@@
More sophisticated numerical analysis shows mathematically more complex correlation functions between forward and backward propagating wave fields with zero average phases, which is shown in the quantitative analysis of the next chapter III. The Schrödinger equation in Eq. (12) once more indicates that spontaneous symmetry breaking isn’t related to the occurrence of a random, but to a pre-defined phase of the quantum field, since the physical constraints in a Bose gas intrinsically pick out a specific zero phase due to quantum coherence at the Boltzmann equilibrium below the critical temperature, while the total distribution of all possible gauge fields is still symmetry-preserving.
@@ II.6 Spontaneous phase gauge symmetry breaking
The concept of spontaneous phase gauge symmetry-breaking explains a certain sudden change in the symmetry properties of the gauge fields that describe a solid below the critical temperature. As an example, one may consider a magnet that changes its magnetization from average zero to a well-defined direction with components of spin up and spin down in the z-direction of the magnet below a specific Curie temperature for a characteristic phase transition. While the experimental observation of symmetry breaking is understood straightforwardly, the formulation of a consistent theory that explains spontaneous symmetry breaking as a physical rather than a purely mathematical process is highly non-trivial.
Figure 3: (Color online) Figure shows the distribution of the (unweighted) condensate wave field modes with the constraint of the Boltzmann equilibrium at the same experimental conditions as in Fig. 2. Instead of the continuously distributed wave field in the complex plane, the wave field shows a star-shaped linear spreading from zero to a finite value in the direction of quantized angles $\phi$.
Any initial state of the formal solution in Eq. (13) with a specific initial phase defined by $\psi_{0}=|\psi|{\rm e}^{i\phi_{0}}$ defines a separate and non-unique solution of the Schrödinger equation. The corresponding gauge fields are symmetric as long as no physical constraint is assumed for the quantum field’s chemical potential and are defined by Glauber coherent states that all preserve the symmetry of the Schrödinger equation. Considering only manifolds of the Boltzmann equilibrium with vanishing chemical potential $\mu=0$ leads to projections onto subspaces of the gauge fields as a consequence of a physical constraint, the Boltzmann equilibrium, i.e., to the distribution $P(\alpha)=\delta(\alpha-\alpha_{0})$, without the requirement of formally defining a certain symmetry-breaking gauge for the wave field. Please note, however, that this does not define a symmetry-breaking process of the quantum system, since the symmetry of the total quantum field is preserved. The projection of the total wave field onto the manifold of the Boltzmann equilibrium finally implies the propagation of a one-directional time, i. e. ”the forward direction of time”. From Eqs. (6) - (8), it is thus likely to understand that physical time passes in only one direction, since at the Boltzmann equilibrium, where disjoint forward and backward propagating components of the quantum field are effectively mixed and equal, the partial waves interfere and lead to a well-defined mathematical form of the quantum field that doesn’t diverge numerically also below the critical temperature for Bose-Einstein condensation.
@@
The breaking of the relative phase gauge symmetry between condensate and non-condensate fields during the phase transition can be theoretically understood as an effect of quantum coherence below the critical temperature that changes the total field from a state of the form $\psi=\psi_{0}\cup\psi_{\perp}$ to a coherent superposition of the components $\psi_{0}$ and $\psi_{\perp}$. Considering a quantum field without local representations of the particle’s wave functions facilitates the formulation of the relative symmetry breaking between the different components (aggregate phases) of the quantum system. As will be illustrated numerically ab initio in the following chapter III, in mathematical terms, symmetry breaking in ideal Bose gases below the critical temperature can be understood by the introduction of a non-local order parameter as defined by a coherent sum of condensate and non-condensate fields,
@@
within the framework of the present number-conserving quantum field theory. Since a total number of $N$ non-interacting atoms is assumed, the total average field $\psi=0$ (the gauge of the order parameter of the total field) vanishes (see e.g. Ref. ref-12 ). This is, in particular, described and accounted for by Eq. (9) which represents the conservation of the total number of particles. From this assumption, one is able to recognize that the condensate and non-condensate aggregate phases of the Bose gas are separated by an angle of $\phi=\pi$, i.e.
@@
hence, interpreting the spontaneously broken gauge symmetry of the order parameter below the critical temperature is possible in terms of the fields in Eq. (17).